OpenStudy (anonymous):

B is between A and C. AB=2x-4 BC=7x+5 AC=82

OpenStudy (anonymous):

What am I trying to find?

OpenStudy (anonymous):

What is the value of x?

OpenStudy (anonymous):

Ah, ok. Here we go! Well because the whole like is AC, then that's the total And AB and BC together, add up the the total 2x - 4 + 7x + 5 = 82 9x + 1 = 82 -1 -1 9x = 82 /9 /9 x = 9.1